# Explainable Anomaly Detection in Images and Videos: A Survey

### v2

@@ IV-C Commmons
Figure 7: Error-based anomaly detection methods work for both image and video data.(a) Schematic diagram of error-based image anomaly detection example adapted from [84]. The error is the distance between the input image and the reconstruction and is minimized during training. The training data are all number 8s, which are the normal samples. In the inference face, the error will be small for 8s but large for other numbers. (b) Schematic diagram of error-based video anomaly detection example with a window length of $7$ adapted from [83]. The consecutive frames can be concatenated in channel dimension and then sent into the network for reconstruction. The example here shows the motion branch approach [83], where the input is raw frames, and the target is the corresponding optical flow maps. The outputs of the network are required to be close to the target optical flows during training. During the training, all samples are normal data containing normal behavior of pedestrians like walking. During the testing, when encountering normal samples (first column: a walking person), the total error can be minimal since each frame’s reconstruction closely approximates the target images. For object anomaly (second column: an unseen object like a car), the reconstruction error of every frame will be large, which shares the same mechanism as error-based image anomaly detection, leading large total error. For motion anomaly (third column: a person catching a dropping bag while looking upwards), the reconstruction error of certain few frames can be small (last three here), but due to the abruptly changing motion, there must be some frames with extraordinarily large error (first four here), leading to a relatively large error in total.
All the error-based image anomaly detection methods can be flexibly adapted to video anomaly detection tasks. The practice is simple: for a video $X\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the video clip length, $C$ denotes the channel number and $H,W$ represents the height and width of the frames respectively. Then we can simply reshape the video tensor by concatenating the channel dimension across the frame level, then we have $X^{\prime}\in\mathbb{R}^{TC\times H\times W}$ and we are able to view it as an image and apply the image anomaly detection method. The rationale that they can still work is that the error-based score usually presents in the form of the subtraction between input and the reconstruction output in some form, so for the object anomaly, the error is certainly discriminative and for the motion anomaly, the error is still discriminative. More concretely, for appearance (object) anomalies in video anomaly detection, including unseen vehicles, skateboards, bicycles, etc., the error score in each frame is discerning, therefore the summation of them across the temporal dimension is still discerning. In this case, the error gap mainly exists at the spatial level. For motion anomalies, despite the fact that looking at each frame individually, the error score might not be that discriminative since the objects may have been witnessed during training (people). Nevertheless, the variation of the pixel values along the time axis has not been seen during the training, so the error gap considering all the consecutive $N$ frames will be large. In this case, the error gap mainly exists at the temporal level. We illustrate such practice and phenomenon in Fig. 7.
It is noteworthy that the error-based score not necessarily take advantage of the reconstruction error. As long as the error-based score accumulates the error over the spatial dimension in some form, it can be easily adapted to video tasks, including reconstruction-error-based score [58, 23, 34, 59], log-likelihood score [59], Dice score [47, 48] and reference discrimination score using GAN [40]. We can also introduce a more refined design along the temporal dimension for better time variation information excavation when adjusting such approaches to video data and we leave this for future work.
Similarly, the attention-based explainable method can be utilized for image anomaly detection. Although the video has a temporal dimension, most methods [61, 23] regard a frame cube as a single object. For example, for a video $X\in\mathbb{R}^{T\times C\times H\times W}$, the sample of a cube has dimension $\mathbb{R}^{C\times H\times T*W}$. In this case, the video sequence can be converted to a large image which actually concatenates several frames. The attention map can thus reflect the predicted anomaly location and visualized weights of each pixel.
Some intrinsic methods [68, 67] can also be transferred to image anomaly detection. The key idea of the intrinsic method is to leverage its own attributes to explain the prediction. For example,  [68] generated three representations of object category, interaction, and size of bounding box and trained a Gaussian Mixture Model (GMM) using these representations. It can explain itself by telling which part is deviated from the normal distribution.  [67] trained three classifiers to classify one’s category, action, and attributes. Although these classifiers are not involved in detecting anomalies, they can show why an object is abnormal afterward. However, as we know, image anomaly detection may not have such attributes as interaction, or motion. Object category usually is not meaningful as the image anomaly detection task has the same object type. Although these video-specific attributes cannot be transferred to images, images themselves have their own attributes, like size, and shape. This additional prior information is helpful to explain the anomaly in a semantic way. Explainable methods in image anomaly detection can definitely adapt a similar procedure to explain the results using image-specific attributes. Besides, some perturbation methods can also adopt such feature embeddings. For example,  [20] perturbed some feature embeddings by the corresponding classifiers.
Foundation-model-based IAD approaches, especially the LLM-based ones can be easily adapted and applied to video data. The difference is that for video anomaly detection, the input visual data has an additional temporal dimension, which can be processed by pooling [85] or learning modules [66, 86] to eliminate the temporal dimension and sent into LLM after concatenation with prompt features. The input token space is inherently capable of handling the temporal relationship of the video frames, making the whole anomaly detection pipeline transparent and interpretable.
@@ V Datasets
In this section, we summarize the most widely used public datasets for 2D anomaly detection, which contributes to the understanding of the explainabilities of the methods applied to them.
V-A Image Anomaly Detection
For OOD detection and semantic anomaly detection works of image anomaly detection, the experiments are usually conducted on image classification datasets by setting some class(es) as normal data and the remainder as abnormal data. The commonly used classification datasets include CIFAR-10 (C10) [24], CIFAR-100 (C100) [24], MNIST (MN) [87], Fashion-MNIST (fMN) [88], Texture [89], LSUN [90], Caltech-101 (Cal) [91], CURE-TSR [92], iSUN [93], SVHN [94] and CelebA [95] etc. The types of explainable IAD methods corresponding to the aforementioned datasets are mainly input-perturbation-based and generative-model-based methods, which concentrate on sample-level discriminativity. So in the following IAD datasets listed, we mainly focus on sensory anomaly detection/localization datasets, which are mostly applicable to attention-based, generative-model-based, and foundation-model-based IAD approaches.
MVTec AD (MVTAD) [25, 96]
presents an extensive dataset designed for unsupervised anomaly detection within natural imagery. This dataset simulates real-world scenarios encountered in industrial inspections and includes 5,354 high-resolution pictures showcasing five distinct textures and ten different objects from various fields. The data set includes 73 different types of anomalies that present as defects or structural changes in objects or their textures. For every image containing a defect, it offers pixel-precise ground truth areas, which enable the assessment of techniques for both anomaly identification and segmentation.
MVTec LOCO AD (MVTLC) [97]
presents a novel dataset aimed at assessing the performance of unsupervised algorithms for localizing anomalies. This dataset comprises 3,644 images across five unique object categories, modeled after real-world industrial inspection settings. The anomalies within the dataset are categorized into structural and logical types. It clearly differentiates between structural anomalies, which are described as physical defects like scratches, dents, or impurities, and logical anomalies, which are characterized by violations of inherent rules, such as an object being in the wrong place or missing entirely.
VisA [98]
is a comprehensive collection of 10,821 images, featuring 9,621 normal and 1,200 anomalous samples across 12 subsets corresponding to different objects, including complex printed circuit boards (PCB) and simpler objects like Capsules, Candles, and various food items. These subsets capture a range of conditions, from multiple instances with varied poses and locations to roughly aligned objects, with anomalies such as surface and structural defects meticulously introduced to simulate realistic conditions. The dataset, captured with a high-resolution RGB sensor, offers both image and pixel-level annotations to support detailed analysis.
Fishyscapes [99]
serves as a public standard for gauging uncertainty in the practical task of semantic segmentation related to urban driving. It assesses the accuracy of pixel-level uncertainty predictions in identifying unusual objects in the vehicle’s path. It is the first public benchmark specifically designed for anomaly detection in a real-world task of semantic segmentation for urban driving scenarios. The benchmark reveals that anomaly detection remains a challenging task, even in common situations. The Fishyscapes dataset is based on data from Cityscapes [100], a well-known benchmark for semantic segmentation in urban driving. The benchmark comprises two main components: Fishyscapes Web, where Cityscapes images are overlaid with objects sourced from the web, and Fishyscapes Lost and Found, which builds on a road hazard dataset collected similarly to Cityscapes.
RoadAnomaly21 [101]
significantly enhances anomaly segmentation in street scenes by offering an evaluation dataset of 100 images, each with detailed pixel-level annotations. This dataset, an expanded and refined version of a prior collection, features a more diverse set of images with improved quality and more precise labeling, including the addition of a void class and 68 new images. It rigorously removes low-quality images and corrects previous labeling errors, ensuring a clear focus on road scenes. Each image in the dataset showcases at least one anomalous object, such as unexpected animals or unidentified vehicles, sourced from a variety of web-based environments to ensure a wide diversity in object sizes, locations, and scene types. To aid in method validation, the dataset also includes 10 extra images with annotations, allowing users to test the compatibility of their approaches with this benchmark.
LAG [102]
, a large-scale glaucoma detection dataset, features $5,824$ fundus images ($2,392$ positive and $3,432$ negative for glaucoma) from Beijing Tongren Hospital, evaluated under the Helsinki Declaration guidelines and exempt from patient consent due to its anonymized nature. Each image has been rigorously diagnosed by specialists using both morphologic and functional analyses, establishing a gold standard for glaucoma detection.
UK Biobank (UKB) [103]
is a vast and comprehensive prospective study initiated between $2006$ and $2010$, enrolling over $500,000$ participants aged 4040404069696969 years, aimed at investigating the genetic and environmental determinants of diseases prevalent in middle and old age. It encompasses a wide array of data including phenotypic and genotypic information, physical measurements, and longitudinal health outcomes, making it a valuable resource for public health research accessible to bona fide researchers globally.
MSLUB [104]
comprises MRI images of $30$ Multiple Sclerosis patients, featuring white-matter lesion segmentations validated through a multi-rater consensus. This dataset, designed for the creation and testing of automated lesion segmentation methods, emphasizes the use of semi-automated tools to reduce variability both within and between different raters, aiming to enhance the accuracy and reliability of MS lesion identification and quantification in clinical settings.
Brain Tumor Segmentation (BRATS) [105]
is designed for the assessment of cutting-edge techniques in segmenting brain tumors using multi-parametric MRI imaging. The collection features pre-operative MRI scans from multiple institutions focused on glioma tumors. It underscores the delineation of different tumor sub-regions and the forecasting of patient survival outcomes, utilizing radiomic characteristics and machine learning techniques.
White Matter Hyperintensities (WMH) [106]
provides a standardized framework for the automatic segmentation of white matter hyperintensities from multi-center, multi-scanner brain MRI data, aiming to facilitate the development and comparison of segmentation methods. It includes $60$ training images with manual segmentations and a test set of $110$ images, evaluated on metrics like Dice similarity coefficient and sensitivity, to validate segmentation algorithms’ performance across different MRI scanners and protocols.
CheXpert [107]
is a large public collection of $224,316$ chest radiographs from $65,240$ patients, labeled for the presence of 14 observations as positive, negative, or uncertain. It includes an innovative approach to capture uncertainties in radiology reports through automatic labeling, facilitating research into the development of algorithms for chest radiograph interpretation that can handle uncertainty effectively.
MedMNIST [108]
is a collection of $10$ pre-processed medical image datasets, designed for classification tasks with lightweight 28x28 images, aiming to facilitate educational purposes, rapid prototyping, and AutoML benchmarks in medical image analysis. It includes diverse data scales and tasks across primary medical image modalities, making it accessible and useful for multi-modal machine learning challenges without requiring extensive background knowledge.
KSDD2 [109]
consists of over $3,000$ high-quality images capturing various types of defects on production items, designed to propel advancements in the detection of surface defects within industrial environments. This dataset, enriched with fine-grained segmentation masks for precise defect localization, represents a significant resource for developing and benchmarking advanced image analysis algorithms for real-world industrial quality control challenges.
Magnetic Tile Defect (MTD) [110]
focuses on the real-time detection of surface defects on magnetic tiles using the MCuePush U-Net model, combining MCue U-Net and Push network for enhanced defect localization. It significantly improves the speed and accuracy of defect detection in industrial settings, reducing inspection time from $0.5$s per image to $0.07$s while offering state-of-the-art performance for identifying defects through deep learning techniques.
SDD [111]
is designed for the segmentation and detection of surface defects, utilizing a deep learning architecture optimized for this task. It is distinguished by its efficiency in learning from a limited number of defective samples, demonstrating its practicality for industrial applications where only a small number of defective examples may be available for training.
BeanTech Anomaly Detection (BTAD) [112]
is a collection of $2,830$ real-world industrial images of three different products, designed for anomaly detection tasks. It features both normal and defective product images, with the defective ones being manually labeled for pixel-wise ground-truth anomaly localization, facilitating the development and evaluation of anomaly detection and localization algorithms
DAGM [113]
was created for the DAGM 2007 competition, focusing on weakly supervised learning for industrial optical inspection. It is designed to challenge and benchmark algorithms in the area of visual defect detection on various manufacturing materials, emphasizing the importance of pattern recognition and machine learning techniques in industrial quality control.
MPDD [114]
addresses the challenge of defect detection in metal fabrication and painting, focusing on the variability and complexity of real-world manufacturing conditions. It aims to improve the performance of semi-supervised and unsupervised anomaly detection methods by providing a dataset that includes a variety of metal parts under different conditions, highlighting the necessity of testing anomaly detection methods in diverse, application-specific scenarios for further development and accuracy enhancement in industrial settings.
V-B Video Anomaly Detection
@@
TABLE V: Comparisons of different Video Anomaly Detection datasets. The UMN dataset does not have official training and testing splits. The UBnormal dataset includes a validation set, which is not displayed here. For setting, “Se” denotes Semi-supervised, “We” denotes Weakly-supervised and “Su” denotes Supervised.
Dataset Year Setting #Scenes # Frames # Anomaly classes Anomaly classes
@@
Subway Entrance [115] 2008 Se 1 144,250 76,453 67,797 5 Loitering, Wrong direction, Illegal exit, Unattended objects, Crowding
@@
USCD Ped 1 [117] 2010 Se 1 14,000 6,800 7,200 5 Biking, Skating, Small carts, People in wheelchair, People walking across a walkway or in the grass that surrounds it
USCD Ped 2 [117] 2010 Se 1 4,560 2,550 2,010 5 Biking, Skating, Small carts, People in wheelchair, People walking across a walkway or in the grass that surrounds it
@@
ShanghaiTech [119] 2017 Se 13 317,398 274,515 42,883 11 Biking, Running, Loitering, Vehicle, Skateboarding, Climbing, Throwing objects, Fighting, Jaywalking, Gathering, Abnormal Object
UCF-Crime [120] 2018 We - 13,741,393 12,631,211 1,110,182 13 Abuse, Arrest, Arson, Assault, Road Accident, Burglary, Explosion, Fighting, Robbery, Shooting, Stealing, Shoplifting, Vandalism

### v4

@@ IV-C Commonalities and Transferable Approaches
Figure 6: Error-based anomaly detection methods work for both image and video data.(a) Schematic diagram of error-based image anomaly detection example adapted from [129]. The error is the distance between the input image and the reconstruction and is minimized during training. The training data are all number 8s, which are the normal samples. In the inference face, the error will be small for 8s but large for other numbers. (b) Schematic diagram of error-based video anomaly detection example with a window length of 77 adapted from [130]. The consecutive frames can be concatenated in channel dimension and then sent into the network for reconstruction. The example here shows the motion branch approach [130], where the input is raw frames, and the target is the corresponding optical flow maps. The outputs of the network are required to be close to the target optical flows during training. During the training, all samples are normal data containing normal behavior of pedestrians like walking. During the testing, when encountering normal samples (first column: a walking person), the total error can be minimal since each frame’s reconstruction closely approximates the target images. For object anomaly (second column: an unseen object like a car), the reconstruction error of every frame will be large, which shares the same mechanism as error-based image anomaly detection, leading large total error. For motion anomaly (third column: a person catching a dropping bag while looking upwards), the reconstruction error of certain few frames can be small (last three here), but due to the abruptly changing motion, there must be some frames with extraordinarily large error (first four here), leading to a relatively large error in total.
Error/Reconstruction-based Methods Many generative models used for IAD (AEs, VAEs, GANs, Flows, Diffusion Models) rely on reconstruction error or likelihood deviation as the anomaly score [54, 83, 82, 60, 84]. The resulting error maps (difference between input and reconstruction) serve as a basic form of explanation by highlighting where the model failed to reconstruct accurately. This principle can be extended to videos, often by treating video clips as sequences of images or concatenating frames channel-wise (Fig. 6). The rationale is that both object anomalies (consistently high error across frames) and motion anomalies (high error in frames depicting abnormal temporal changes) will result in larger overall reconstruction errors compared to normal sequences. While adaptable, explaining why the reconstruction failed (e.g., due to unseen object shape vs. unexpected motion) might require further analysis beyond the raw error map.
Attention Mechanisms Attention maps, whether post-hoc (like Grad-CAM [103]) or learned [48], highlight salient regions contributing to the model’s decision. This concept is applicable to both IAD [37, 50] and VAD [85]. For videos, attention can be computed frame-wise or across spatio-temporal dimensions. The resulting attention maps offer spatial localization explanations. However, as noted [114], attention might simply highlight foreground objects rather than the anomaly itself, limiting its explanatory power. Evaluating the faithfulness of attention maps (i.e., whether they truly reflect the model’s reasoning) is crucial [131].
Foundation Models (VLMs/LLMs) Approaches using large pre-trained models like CLIP [108] or GPT-4V [110] offer promising transferability [66, 67, 68, 69, 71, 72]. By projecting visual features into a shared space with text embeddings, these models can leverage natural language for explaining anomalies. The same fundamental architecture (visual encoder + language model) can process images or adapted video representations (e.g., using frame pooling [132] or temporal modeling modules [90, 133]). The explanation is generated via prompting the language model (e.g., ”Describe the anomaly in this image/video.”). The key challenge lies in ensuring the generated textual explanations are accurate, relevant, and truly reflect the visual anomaly, rather than being plausible but incorrect hallucinations.
Shapley Values / Feature Attribution Techniques like SHAP [125, 123, 124, 126], which attribute prediction changes to input features, are theoretically applicable to both images (attributing to pixels/patches) and videos (attributing to spatio-temporal regions). They provide fine-grained, mathematically grounded explanations. However, their computational cost can be prohibitive for high-dimensional video data, and interpreting pixel-level attributions for complex temporal events can still be challenging.
Challenges in Explainable Visual Anomaly Detection Across both modalities, key challenges remain. Defining and quantifying the quality of explanations is difficult. Standard AD metrics (AUC, IoU) measure detection/localization performance, not explanatory power. Evaluating aspects like faithfulness (does the explanation accurately reflect the model’s reasoning?), stability (does the explanation change significantly with small input perturbations?), and effectiveness (does the explanation help the user understand or debug the model?) requires specialized metrics and user studies [131, 17]. Furthermore, relating low-level explanations (e.g., attention maps, pixel attributions) to high-level semantic concepts understandable by humans is often non-trivial. Bridging this gap, perhaps by combining different explanation techniques or leveraging foundation models, is an important direction.
@@ V Datasets
In this section, we summarize commonly used public datasets for 2D visual anomaly detection. Understanding the characteristics of these datasets is crucial for evaluating anomaly detection performance and the effectiveness of explainability methods, particularly how well explanations capture the nature of anomalies present in different domains.
V-A Image Anomaly Detection Datasets
A variety of datasets are used for IAD, ranging from standard image classification benchmarks adapted for OOD/semantic anomaly detection to specialized datasets focusing on industrial defects or medical anomalies.
General/OOD Datasets
For OOD detection and semantic anomaly detection (where one or more classes are treated as normal and others as anomalous), standard image classification datasets are often repurposed. Examples include CIFAR-10 (C10) [38], CIFAR-100 (C100) [38], MNIST (MN) [134], Fashion-MNIST (fMN) [135], ImageNet [116], Texture [136], LSUN [137], Caltech-101 (Cal) [138], CURE-TSR [139], iSUN [140], SVHN [141], and CelebA [142]. Explainable methods applied here often focus on sample-level discriminativity (e.g., input perturbation, likelihood scores) to explain why an image is classified as out-of-distribution or belonging to an unseen class. The nature of the anomaly is primarily semantic or a domain shift.
Industrial Defect Datasets
These datasets simulate real-world industrial inspection scenarios and are crucial for evaluating methods aimed at localizing and explaining physical defects. Pixel-level ground truth is often available, facilitating the evaluation of localization-based explanations (e.g., attention maps, reconstruction error maps).
  • MVTec AD (MVTAD) [39, 12] A widely used benchmark with 5,354 high-resolution images across 15 texture and object categories. Features 73 types of defects (structural changes, missing parts, contamination). Pixel-precise annotations enable evaluation of detection and segmentation/localization explanations.
  • VisA [143] 10,821 images (9,621 normal, 1,200 anomalous) across 12 object types (e.g., PCBs, capsules). Includes surface and structural defects with image and pixel-level annotations.
  • MVTec LOCO AD (MVTLC) [144] Focuses on localization, distinguishing between structural anomalies (physical defects) and logical anomalies (rule violations, e.g., misplaced/missing objects). Contains 3,644 images across 5 categories. Evaluating explanations for logical anomalies poses unique challenges.
  • KSDD2 [145] Over 3,000 images with fine-grained segmentation masks for surface defects on production items.
  • Magnetic Tile Defect (MTD) [146] Focuses on real-time defect detection on magnetic tiles.
  • SDD [147] Designed for segmenting surface defects, particularly with limited defective samples.
  • BeanTech Anomaly Detection (BTAD) [148] 2,830 images of 3 industrial products with pixel-wise anomaly annotations.
  • DAGM [149] From a 2007 competition, focuses on weakly supervised defect detection.
  • MPDD [150] Metal parts with defects under various manufacturing conditions.
Explaining anomalies in these datasets typically involves localizing the defect (e.g., via heatmaps) and potentially classifying the defect type. Logical anomalies in MVTLC require explanations that go beyond local appearance, potentially involving reasoning about expected object configurations.
Street Scene / Autonomous Driving Datasets
These datasets focus on identifying unexpected objects or hazards in driving scenarios.
  • Fishyscapes [151] Based on Cityscapes [152], evaluates pixel-level uncertainty for detecting unusual objects (OOD objects overlaid or from a road hazard dataset). Explainability relates to how well uncertainty estimates correspond to semantic anomalies.
  • RoadAnomaly21 [153] Focuses on anomaly segmentation with 100 street scene images with pixel-level annotations for anomalous objects (e.g., unexpected animals, vehicles).
Explanations often involve segmenting the anomalous object and identifying it as out-of-place in the driving context.
Medical Imaging Datasets
Anomaly detection in medical images is critical, and explainability is paramount for clinical trust.
  • LAG [154] 5,824 fundus images for glaucoma detection (anomaly = glaucoma presence).
  • UK Biobank (UKB) [155]: Large-scale resource including various imaging modalities; can be used for diverse anomaly detection tasks.
  • MSLUB [156] MRI images with multiple sclerosis lesion segmentations.
  • Brain Tumor Segmentation (BRATS) [157]: MRI scans for glioma tumor segmentation.
  • BrainMRI [158]: Magnetic resonance images for brain tumor detection, including various tumor types and normal brain scans, processed to evaluate automated detection methods.
  • White Matter Hyperintensities (WMH) [159] Brain MRI data for segmenting white matter hyperintensities.
  • CheXpert [160] Large collection of chest radiographs labeled for 14 observations.
  • MedMNIST [161] Collection of 10 pre-processed medical image datasets for various classification tasks.
Explanations typically involve localizing the pathological region (lesion, tumor) and often relating it to specific medical conditions. The need for high fidelity and clinical relevance makes explainability particularly challenging and important.
A summary of representative IAD datasets highlighting their relevance to explainability is presented in Tab. III.
TABLE III: Summary of representative Image Anomaly Detection (IAD) datasets relevant to explainability. GT = Ground Truth.
Dataset Domain # Images Anomaly Type Pixel GT Explainability Relevance
MVTec AD [39] Industrial 5,354 Structural Defects Yes Localization heatmaps, defect classification
VisA [143] Industrial 10,821 Surface/Structural Defects Yes Localization heatmaps, defect classification
MVTec LOCO AD [144] Industrial 3,644 Structural & Logical Defects Yes Localization, explaining rule violations (logical)
Fishyscapes [151] Driving (Based on Cityscapes) OOD Objects Yes Uncertainty maps as explanation, semantic segmentation
RoadAnomaly21 [153] Driving 100 (+10) OOD Objects Yes Anomaly segmentation maps
BRATS [157] Medical (MRI) Varies Brain Tumors Yes Lesion segmentation, clinical relevance
CheXpert [160] Medical (X-ray) 224,316 Pathological Observations No (Image-level labels) Explaining image-level diagnosis (e.g., via attention)
CIFAR-10/100 [38] General 60k/60k Semantic/OOD No Sample-level explanations (perturbation, likelihood)
V-B Video Anomaly Detection Datasets
Video datasets present unique challenges due to the temporal dimension, involving anomalies in actions, interactions, or events unfolding over time.
@@ V-B Video Anomaly Detection Datasets
TABLE IV: Comparisons of different Video Anomaly Detection datasets. The UMN dataset does not have official training and testing splits. The UBnormal dataset includes a validation set, which is not displayed here. For setting, “Se” denotes Semi-supervised, “We” denotes Weakly-supervised and “Su” denotes Supervised.
Dataset Year Setting #Scenes # Frames
# Anomaly
classes
Anomaly classes
@@
Subway Entrance [162] 2008 Se 1 144,250 76,453 67,797 5
Loitering, Wrong direction, Illegal exit,
Unattended objects, Crowding
@@
USCD Ped 1 [164] 2010 Se 1 14,000 6,800 7,200 5
Biking, Skating, Small carts, People in wheelchair,
People walking across a walkway or in the grass that surrounds it
USCD Ped 2 [164] 2010 Se 1 4,560 2,550 2,010 5
Biking, Skating, Small carts, People in wheelchair,
People walking across a walkway or in the grass that surrounds it
@@
ShanghaiTech [166] 2017 Se 13 317,398 274,515 42,883 11
Biking, Running, Loitering, Vehicle,
Skateboarding, Climbing, Throwing objects, Fighting,
Jaywalking, Gathering, Abnormal Object
UCF-Crime [167] 2018 We - 13,741,393 12,631,211 1,110,182 13
Abuse, Arrest, Arson, Assault, Road Accident,
Burglary, Explosion, Fighting, Robbery,
Shooting, Stealing, Shoplifting, Vandalism